# A Reproducible Framework for Neural Topic Modeling in Focus Group Analysis

### v1

@@ 3.4 Model Selection and Validation
Finally, to assess the coherence and interpretability of the merged topics, we conducted a human validation study. Three independent annotators, with expertise in public health and qualitative analysis, rated the coherence of each topic on a 5-point Likert scale (1=Incoherent, 5=Highly Coherent) and assigned them interpretable labels. Each topic was represented with a set of randomly selected representative documents. Raters were blind to topic labels and first evaluated coherence before collaboratively assigning interpretable labels. Inter-annotator agreement was assessed using weighted Cohen’s Kappa[mchugh_interrater_2012], which accounts for the ordinal nature of the ratings, and the Intraclass Correlation Coefficient (ICC(2,1)), [shrout_intraclass_1979] which quantifies overall reliability across all three raters.
@@ 3.5 LDA Baseline Comparison
To contextualize the performance of our BERTopic approach, we implemented Latent Dirichlet Allocation as a comparative baseline. Following current practice in applied topic modeling research [hagg2022], we tuned only the number of topics while maintaining default settings for other hyperparameters. Hagg et al.’s scoping review found indeed that most applications leave parameters at defaults, making this approach representative of typical usage.
We performed grid search across the number of topics $\in\{5,7,9,10,12,15\}$, selected to span coarse to fine-grained granularity appropriate for our corpus size. Each configuration was evaluated using the same $C_{v}$ coherence scoring framework applied to BERTopic models. The optimal configuration was selected based on maximum coherence score.
@@ 4 Results
We present results from our systematic hyperparameter exploration, baseline LDA comparison, stability analysis, performance assessment, model selection, and human validation. Our evaluation spanned 27 BERTopic configurations and one LDA baseline.
@@ 4.1 Stability Analysis
We evaluated topic model stability across 27 hyperparameter combinations using bootstrap resampling with 30 replicates. Figure 2 presents stability metrics (NMI and ARI) across 27 hyperparameter combinations.
The two metrics reveal contrasting patterns across the parameter space. NMI values range from 0.376 to 0.700, with highest stability observed for min_cluster_size = 5 configurations (0.624–0.700). Stability decreases substantially for larger cluster sizes: min_cluster_size = 10 achieves 0.480–0.555, while min_cluster_size = 15 shows 0.376–0.439.
ARI values are systematically lower (0.094–0.205) and exhibit the opposite trend. Configurations with min_cluster_size = 15 achieve the highest ARI scores (0.159–0.205), while min_cluster_size = 5 shows the lowest values (0.094–0.171). This divergence reflects fundamental differences in what these metrics quantify: NMI measures information structure preservation and remains stable when fine-grained clusters consistently capture semantic distinctions, while ARI measures exact pairwise assignment agreement and is sensitive to boundary variations in small clusters where minor document shifts represent large proportional changes.
Within each min_cluster_size group, both metrics consistently show that n_neighbors = 10 produces the most stable results, while n_components has more modest effects on stability.
Figure 2: Stability metrics (NMI and ARI) across 27 hyperparameter configurations. NMI favors smaller cluster sizes while ARI shows opposite pattern, reflecting different sensitivities to cluster size.
@@ 4.2 Performance Assessment
Figure 3 shows performance metrics across the entire parameter space. Coherence scores vary substantially across configurations, ranging from 0.385 to 0.589. The highest coherence is achieved by a min_cluster_size = 15 configuration (0.589), while min_cluster_size = 5 configurations show more modest coherence values (0.385–0.427).
The number of discovered topics varies dramatically with min_cluster_size: configurations with min_cluster_size = 5 produce 36–52 topics, min_cluster_size = 10 yields 12–25 topics, and min_cluster_size = 15 generates 2–14 topics. Outlier fractions range from 5.8% to 44.1%, with generally higher outlier rates for smaller cluster size thresholds. Notably, the configuration with highest coherence (min_cluster_size = 15, n_neighbors = 30, n_components = 10) produces only 2 topics, indicating potential over-aggregation that may limit analytical utility.
Figure 3: Topic quality metrics across parameter space. Coherence scores range from 0.385-0.589, with number of topics varying from 2-52 depending on min_cluster_size.
@@ 4.3 Model Selection
Based on the stability and performance assessments, we identified two competing approaches that represent different optimization strategies:
Contender A: Stability-Optimized Approach The most stable configuration (min_cluster_size = 5, n_neighbors = 10, n_components = 5) achieved NMI = 0.695 and ARI = 0.171 and initially produced 49 topics with coherence = 0.392 and 23.6% outliers. Using hierarchical clustering analysis (Figure 4(a)), we merged related topics creating 9 semantically coherent super-topics, and improving coherence to 0.558 while maintaining 23.6% outliers.
(a) Hierarchical clustering dendrogram showing relationships among the 49 initial fine-grained topics.
(b) Composition of the 9 merged topics showing document distribution and initial fragmentation. Bar height represents total documents per topic (shown at top). White lines separate constituent subtopics, with numbers indicating document counts.
Figure 4: Hierarchical topic merging process and resulting composition. (a) The dendrogram reveals natural groupings among the 49 initial fine-grained topics extracted by BERTopic, with colored branches delineating the 9 merged topics. (b) The composition chart quantifies both the importance (bar height) and initial fragmentation (number of subtopics) of each merged topic
Contender B: Direct Coherence-Optimized Approach Among configurations producing a reasonable number of topics (>5), the highest coherence was achieved by min_cluster_size = 15, n_neighbors = 15, n_components = 10, which directly produced 9 topics with coherence = 0.539 and 26.8% outliers.
Final Selection We selected Contender A as our final model based on superior coherence performance (0.558 vs. 0.539) and lower outlier fraction (23.6% vs. 26.8%).
@@ 4.4 LDA Baseline Comparison
Systematic grid search across LDA number of topics (5-15 topics) identified 7 topics as optimal. Table 2 compares the optimized LDA baseline with our final BERTopic model.
@@
LDA (optimized) 7 0.546 N/A
BERTopic (merged) 9 0.558 23.6
Agreement: NMI = 0.122, ARI = 0.080
Table 2: Comparison of optimized LDA baseline and final hierarchical BERTopic model. Despite comparable coherence, minimal agreement (NMI, ARI) indicates fundamentally different topic structures.
The minimal agreement between models (NMI = 0.122, ARI = 0.080) indicates that despite comparable coherence values, LDA and BERTopic discover fundamentally different thematic structures. This divergence likely reflects architectural differences between LDA’s bag-of-words assumptions and BERTopic’s transformer-based contextual embeddings.
@@ 4.5 Final Topic Structure and Human Validation
Human validation by three domain experts confirmed that the 9 merged topics were thematically coherent and interpretable. Inter-annotator agreement was robust, with a mean pairwise weighted Cohen’s Kappa of 0.578 (moderate agreement) and a strong mean Spearman’s rank correlation ($\rho$ = 0.725). The Intraclass Correlation Coefficient (ICC) for absolute agreement was 0.79 (ICC(2,1)), indicating very good reliability per [koo_guideline_2016]. Notably, the annotators’ ratings were in perfect agreement within a one-point tolerance (100.0%), underscoring a high level of consensus on topic coherence.
Table 3 presents the coherence ratings for each topic, along with consensus labels assigned by the annotators. The mean coherence scores ranged from 2.00 to 5.00. Six of the nine topics received high ratings (mean > 4.0), including those related to Historical trust in national vaccine programs and Polarized attitudes towards official public health sources, both achieving perfect scores of 5.00. Conversely, the lowest-rated topic (mean = 2.00) concerned Prevention practices and barriers to care, suggesting that its representative sentences were perceived as less thematically unified. There was no significant correlation between a topic’s size (number of sentences) and its coherence score ($\rho$ = -0.213, p = 0.583), indicating that the perceived quality of a topic was independent of its prevalence in the discourse.
@@
Topic 0 4 4 3 3.67 COVID-19 vaccine side effects and trust erosion
Topic 1 3 4 3 3.33 Conservative norms and sexual health education
Topic 2 5 5 5 5.00 Historical trust in national vaccine programs
Topic 3 4 5 4 4.33 Parental concerns about HPV vaccination
Topic 4 5 5 4 4.67 Information sources and communication channels
Topic 5 4 4 5 4.33 Personal experiences with cancer and social impact
Topic 6 2 2 2 2.00 Prevention practices and barriers to care
Topic 7 5 5 5 5.00 Polarized attitudes towards official public health sources
Topic 8 4 5 4 4.33 General disease awareness and knowledge gaps
Inter-rater reliability: Mean weighted Cohen’s $\kappa$ = 0.578; Mean Spearman’s $\rho$ = 0.725; ICC(2,1) = 0.79
Table 3: Human validation results for the 9 merged topics. Coherence scores range from 1 (incoherent) to 5 (highly coherent). Representative utterances for each topic are provided in Appendix A.
The 9 emerging themes range from general awareness and personal experiences with cancer to specific concerns about vaccine side effects and preferred communication channels. Figure 5 visualizes the semantic organization of the topics in two-dimensional embedding space, revealing distinct thematic clusters and their spatial relationships. While this paper focuses on methodological validation, the substantive themes identified merit deeper analysis to understand their public health implications. A forthcoming companion paper will provide detailed thematic analysis and discuss policy recommendations based on these findings.
Figure 5: Two-dimensional visualization of document embeddings colored by final merged topic assignments. Each point represents a document (utterance) from the focus group transcripts. Colored regions show the 9 merged topics, with spatial proximity indicating semantic similarity.
@@ 5 Discussion
5.1 Bootstrap Stability Analysis and the NMI-ARI Divergence
Our implementation of bootstrap stability analysis builds on Greene et al. [greene2014]’s recommendations of performing bootstrap resampling for assessing topic model robustness. A critical empirical finding concerns the divergent behavior of NMI and ARI metrics across the parameter space. While NMI values ranged from 0.376 to 0.700 with highest stability at smaller cluster sizes (min_cluster_size = 5), ARI showed the opposite pattern with values of 0.094 to 0.205 and peak performance at larger cluster sizes (min_cluster_size = 15).
@@ 5.1 Bootstrap Stability Analysis and the NMI-ARI Divergence
This divergence reflects the fundamental differences in what these metrics quantify. ARI (Adjusted Rand Index) measures exact agreement in pairwise cluster assignments [hubert1985]. It is highly sensitive to boundary shifts: when documents near cluster boundaries are reassigned across bootstrap samples, this creates pairwise disagreements that lower ARI scores. This sensitivity is particularly pronounced for small clusters, where even minor document reassignments represent large proportional changes to cluster composition. In contrast, NMI (Normalized Mutual Information) measures whether clustering solutions preserve the same underlying information structure based on entropy [strehl2002]. NMI remains relatively stable when fine-grained semantic distinctions are consistently recovered, even if exact cluster boundaries shift somewhat across bootstrap samples.
Recent systematic evaluations of topic modeling algorithms have demonstrated that ARI and NMI can favor different model configurations depending on data characteristics [rudiger_topic_2022], confirming that neither metric is universally optimal. For our exploratory analysis of focus group discussions, where the goal is to discover the full thematic landscape including both dominant and peripheral themes, we argue that NMI provides a more appropriate stability criterion. Fine-grained, information-rich models better capture the nuanced structure of focus group discourse, including smaller but important sub-themes that inform health communication interventions.
5.2 Navigating the Stability-Coherence Tradeoff
The stability-coherence tradeoff we observed, with smaller cluster sizes yielding higher stability (NMI = 0.695) but lower coherence (0.392), while larger cluster sizes produced higher coherence (0.589) but lower stability (NMI = 0.376), is an example of the multi-dimensional optimization challenge discussed by Hosseiny and Baumer [hosseiny2023], who emphasized that no single optimal configuration exists for all contexts. Our hierarchical merging strategy offers a practical solution: by extracting 49 fine-grained topics with high stability then consolidating to 9 super-topics, we achieved coherence (0.558) superior to direct extraction of 9 topics (0.539) while maintaining stability benefits.
The hierarchical approach also addresses a practical challenge: fine-grained models capture nuanced distinctions but may overwhelm readers, while coarse models sacrifice detail for simplicity. By explicitly modeling topic hierarchies, researchers can present high-level summaries while preserving access to fine-grained distinctions for detailed analysis, which is particularly valuable for focus group research where different stakeholders require different levels of granularity.

### v2

@@ 3.5 Human Validation
To assess the coherence and interpretability of our final topic model, we conducted a validation study with three independent annotators having expertise in public health and qualitative analysis. Each annotator rated the coherence of each topic on a 5-point Likert scale (1=Incoherent, 5=Highly Coherent) based on 10 representative documents randomly selected from each topic. Raters were blind to topic labels and model-generated keywords, and first evaluated coherence independently before collaboratively assigning interpretable labels to each topic.
Inter-annotator agreement was assessed using weighted Cohen’s Kappa[mchugh_interrater_2012], which accounts for the ordinal nature of ratings, and the Intraclass Correlation Coefficient (ICC(2,1))[shrout_intraclass_1979], which quantifies overall reliability across all three raters.
@@ 3.6 LDA Baseline Comparison
To contextualize the performance of our BERTopic approach, we implemented Latent Dirichlet Allocation as a comparative baseline. Following current practice in applied topic modeling research [hagg2022], we tuned only the number of topics while maintaining default settings for other hyperparameters. Hagg et al.’s scoping review found that most applications leave parameters at defaults, making this approach representative of typical usage.
We performed grid search across the number of topics $\in\{3,\cdots,15\}$, selected to span coarse to fine-grained granularity appropriate for our corpus size. Each configuration was evaluated using the same $C_{v}$ coherence scoring framework applied to BERTopic models. The optimal LDA configuration was selected based on maximum coherence score.
@@ 4.1 Stability Analysis
We evaluated topic model stability across 27 hyperparameter combinations using bootstrap resampling with 30 replicates per configuration. This analysis showed NMI values ranging from 0.388 to 0.703 (median: 0.558), while ARI values ranged from 0.107 to 0.202 (median: 0.156). Figure 2 (top panels) presents these stability metrics across the parameter space.
The two metrics reveal contrasting patterns. NMI values are highest for min_cluster_size = 5 configurations (0.63-0.70), decreasing substantially for larger cluster sizes: min_cluster_size = 10 achieves 0.48-0.56, while min_cluster_size = 15 shows 0.39-0.44. ARI values are systematically lower and exhibit the opposite trend: configurations with min_cluster_size = 15 achieve the highest ARI scores (0.15-0.20), while min_cluster_size = 5 shows the lowest values (0.11-0.16).
This divergence reflects fundamental differences in what these metrics quantify. NMI measures information structure preservation and remains stable when fine-grained clusters consistently capture semantic distinctions, even when exact boundaries shift across bootstrap samples. ARI measures exact pairwise assignment agreement and is highly sensitive to boundary variations, particularly in small clusters where minor document shifts represent large proportional changes. The negative correlation between NMI and ARI (r = -0.691) further illustrates their divergent behavior.
Critically, we observed a strong negative correlation between NMI and model coherence (r = -0.911, p < 0.001). Configurations with highest NMI (0.63-0.70) produced the least coherent topics (coherence = 0.39-0.43), while configurations with lower NMI (0.39-0.44) achieved substantially higher coherence (0.52-0.59). In contrast, ARI demonstrated a moderate positive correlation with coherence (r = 0.639, p < 0.001), though the universally low ARI range limits its practical utility for model selection. These patterns indicate that stability as measured by NMI and semantic quality as measured by coherence capture fundamentally different constructs, with implications for model selection strategies.
@@ 4.2 Performance Assessment
Figure 2 (middle and bottom panels) displays performance metrics across the entire parameter space. Coherence scores vary substantially, ranging from 0.387 to 0.587 (mean = 0.478, SD = 0.060). The highest coherence is achieved by coarser configurations with min_cluster_size = 15 (0.50-0.59), while min_cluster_size = 5 configurations show lower coherence (0.39-0.43) despite higher stability by NMI.
Figure 2: Quality metrics across the complete hyperparameter space. Six metrics are displayed as heatmaps across three minimum cluster sizes (columns). Each heatmap shows UMAP n_neighbors (rows: 10, 15, 30) by UMAP n_components (columns: 5, 10, 50). Top panels show bootstrap stability (NMI and ARI), middle panels show topic quality (coherence and number of topics), and bottom panels show coverage metrics (outlier fraction and main topic proportion).
The number of discovered topics varies dramatically with min_cluster_size: configurations with min_cluster_size = 5 produce 34-55 topics, min_cluster_size = 10 yields 10-25 topics, and min_cluster_size = 15 generates 2-12 topics. Outlier fractions range from 0.7% to 38.8%, with generally higher outlier rates for finer-grained configurations. Main topic proportions range from 4% to 66%, indicating substantial variation in topic size distributions.
@@ 4.3 Model Selection
Table 2 presents the top 10 configurations ranked by coherence score.
Rank $n$ nc mcs Topics Coherence NMI ARI Outliers
1 10 10 15 2 0.587 0.435 0.167 0.007
2* 15 10 15 7 0.573 0.388 0.164 0.233
3 10 50 15 8 0.549 0.420 0.165 0.249
4 30 10 15 10 0.523 0.420 0.195 0.356
5 30 5 10 14 0.521 0.507 0.146 0.388
6 15 10 10 11 0.519 0.531 0.151 0.286
7 30 50 15 10 0.516 0.419 0.189 0.371
8 15 5 15 10 0.515 0.416 0.153 0.236
9 30 5 15 9 0.511 0.420 0.202 0.335
10 10 5 15 12 0.504 0.431 0.163 0.367
Table 2: Top 10 configurations ranked by coherence score. Configurations are identified by base hyperparameters (n_neighbors ($n$), n_components(nc), and min_cluster_size (mcs)). The selected configuration (rank 2, marked with *) balances high coherence with reasonable granularity.
Based on our multicriteria selection approach, we selected the configuration ranking second by coherence: n_neighbors = 15, n_components = 10, min_cluster_size = 15, which discovered 7 non-outlier topics. This configuration achieved coherence of 0.573 with bootstrap stability of NMI = 0.388 and ARI = 0.164, assigning 23.3% of documents to outliers with the largest topic containing 32.5% of non-outlier documents.
The selection rationale balanced multiple criteria: while rank 1 achieved marginally higher coherence (0.587), its collapse to only 2 topics was deemed too coarse for meaningful exploration of vaccine perceptions. Our selected model provides sufficient granularity (7 topics) to capture distinct thematic dimensions while maintaining strong coherence. The moderate stability metrics reflect the broader pattern that high-coherence models show lower NMI due to coarser clustering, consistent with the strong negative correlation documented above.
@@ 4.4 LDA Baseline Comparison
We compared our BERTopic model against Latent Dirichlet Allocation using the same coherence-based evaluation framework with a grid search across 3 to 15 topics, and evaluated each configuration using Cv coherence.
LDA coherence scores ranged from 0.366 (15 topics) to 0.486 (3 topics), with performance generally declining as granularity increased. The optimal configuration identified 3 topics with coherence of 0.486. Table 3 compares the optimized LDA baseline with our BERTopic model.
@@
LDA (optimized) 3 0.486 N/A
BERTopic (selected) 7 0.573 23.3
Improvement: +18%
Table 3: Comparison of optimized LDA baseline and final BERTopic model. BERTopic achieves 18% higher coherence while discovering more than twice as many interpretable topics.
Our BERTopic model substantially outperformed the baseline, achieving 18% higher coherence (0.573 vs. 0.486) while discovering more than twice as many interpretable topics (7 vs. 3). This performance gap likely reflects architectural differences: LDA’s bag-of-words assumptions treat documents as unordered token collections, while BERTopic’s transformer-based embeddings capture contextual semantics. For conversational data where meaning depends on discourse context (such as “je n’ai pas confiance” referencing different trust domains across utterances) contextualized representations appear better suited to capturing semantic nuance. The LDA baseline’s optimal configuration at only 3 topics suggests that without contextualized embeddings, the model struggles to differentiate finer semantic distinctions in French-language focus group discourse.
@@ 4.5 Final Topic Structure and Human Validation
Human validation by three independent domain experts confirmed topic coherence and interpretability. The seven emerging themes as labeled by the annotators were: General health knowledge and taboos (Topic 0, 32%), Vaccination importance and institutional processes (Topic 1, 23%), Trusted information sources (Topic 2, 9%), Personal vaccination experiences (Topic 3, 5%), Public versus private healthcare debate (Topic 4, 3%), COVID-19 vaccine adverse effects (Topic 5, 3%), and Institutional trust erosion post-pandemic (Topic 6, 1%).
Inter-annotator agreement was robust: ICC(2,1) = 0.700 indicates very good reliability per  [koo_guideline_2016], while mean weighted Cohen’s kappa of 0.678 demonstrates substantial agreement [landis1977]. Annotators achieved perfect agreement within one-point tolerance for 85.7% of topics (6/7).
Table 4 presents coherence ratings for each topic. Five of seven topics received mean ratings $\geq$ 4.0, indicating good-to-excellent coherence. Ratings ranged from 1.67 to 4.67 (overall mean = 3.81). Topic 0 received the lowest rating (mean = 1.67), reflecting its broad, heterogeneous content spanning multiple health domains. Topics 1, 2, 3, 5, and 6 achieved mean ratings between 4.0 and 4.67, confirming strong semantic coherence. Topic 4 received moderate ratings (mean = 3.33). Topic size showed no correlation with coherence ratings ($\rho$ = -0.213, p = 0.583), confirming evaluations reflected semantic quality rather than prevalence.
@@
0 2 2 1 1.67 General health knowledge and taboos
1 5 5 3 4.33 Vaccination importance and processes
2 5 5 4 4.67 Trusted information sources
3 4 5 4 4.33 Personal vaccination experiences
4 3 4 3 3.33 Public vs. private healthcare debate
5 4 5 4 4.33 COVID-19 vaccine adverse effects
6 4 4 4 4.00 Erosion of institutional trust post-COVID
Inter-rater reliability: Weighted Cohen’s $\kappa$ = 0.678; ICC(2,1) = 0.700
Table 4: Human validation results. Coherence scores range from 1 (incoherent) to 5 (highly coherent).
Figure 3 visualizes the semantic organization of the topics in two-dimensional embedding space, revealing distinct thematic clusters and their spatial relationships. While this paper focuses on methodological validation, the substantive themes identified merit deeper analysis to understand their public health implications. A forthcoming companion paper will provide detailed thematic analysis and discuss policy recommendations based on these findings.
Figure 3: Two-dimensional visualization of document embeddings colored by final merged topic assignments. Each point represents a document (utterance) from the focus group transcripts. Colored regions show the 7 merged topics, with spatial proximity indicating semantic similarity.
@@ 5 Discussion
5.1 Bootstrap Stability: A Multifaceted Construct
Our bootstrap stability analysis revealed that partition similarity metrics capture fundamentally different aspects of clustering agreement. NMI and ARI exhibited a strong negative correlation (r = -0.691), systematically disagreeing on which configurations represent stable models. NMI favored fine-grained configurations (min_cluster_size = 5) producing 34-55 topics with high information structure preservation (NMI = 0.63-0.70), while ARI favored coarser configurations (min_cluster_size = 15) with fewer topics where exact pairwise assignments remained more consistent (ARI = 0.15-0.20).
@@ 5.1 Bootstrap Stability: A Multifaceted Construct
This divergence reflects fundamental differences in what these metrics quantify. NMI measures whether clustering solutions preserve the same information structure based on entropy, remaining stable when fine-grained semantic distinctions are consistently recovered across bootstrap samples even if exact cluster boundaries shift [strehl2002]. ARI measures exact pairwise assignment agreement, highly sensitive to boundary variations particularly in small clusters where minor document reassignments represent large proportional changes [hubert1985]. Recent systematic evaluations have confirmed that NMI and ARI can produce different algorithm rankings depending on corpus characteristics [rudiger_topic_2022], indicating that "stability" is not a monolithic construct but depends on which aspect of partition agreement is prioritized.
Moreover, these stability metrics showed different relationships with topic interpretability. NMI exhibited a strong negative correlation with coherence (r = -0.911), while ARI showed moderate positive correlation (r = 0.639) though with limited discriminative power due to universally low values. This metric-dependent relationship demonstrates that bootstrap stability and semantic quality represent distinct dimensions that may conflict rather than align. These findings suggest that stability assessment should employ multiple metrics to capture different facets of partition robustness, and that stability measures alone provide insufficient guidance for model selection in exploratory research contexts where interpretability is paramount.
5.2 Multi-Criteria Model Selection and Validation
The competing objectives revealed by our systematic evaluation (possible stability-coherence tradeoff, varying topic granularity, and different coverage patterns) demonstrate why rigid optimization strategies prove insufficient in practice. While the highest-coherence configuration collapsed to only 2 topics, our selection of 7 topics reflects a judgment that thematic resolution matters alongside semantic quality. This decision acknowledges that optimal configurations depend on research goals: classification tasks might prioritize stability for reproducibility, while exploratory analysis requires balancing interpretability with analytical utility.
Human validation by three domain experts confirmed that coherence-prioritized selection produced interpretable topics, with robust inter-annotator agreement (ICC = 0.700, $\kappa$ = 0.678) and the majority of topics receiving high ratings. Importantly, ratings reflected semantic quality rather than topic prevalence, confirming that computational coherence metrics successfully predict expert judgments of interpretability in our application domain.
The alignment between automated Cv scores and human assessments addresses longstanding concerns that algorithmic optimization may produce statistically optimal but substantively incoherent topics [chang2009]. Our multi-criteria framework provides a transparent, defensible alternative to single-metric optimization or default configurations.